OpenStudy (anonymous):

Without drawing the graph of the equation, determine how many points the parabola has in common with the x-axis and whether its vertex lies above, on, or below the x-axis. y = x2 – 12x + 12 A. 1 point in common; vertex on x-axis B. no points in common; vertex above x-axis C. 2 points in common; vertex below x-axis D. 2 points in common; vertex above x-axis

OpenStudy (anonymous):

@AccessDenied can u help me plz?

OpenStudy (anonymous):

@hartnn can u help?

hartnn (hartnn):

do you know how to find the discrminant of a quadratic expression ?

hartnn (hartnn):

*discriminant

OpenStudy (anonymous):

@hartnn b^2(a*c*4)/b-2 right?

hartnn (hartnn):

not actually, for \(ax^2+bx+c=0\) the discriminant D \(\large D =b^2-4ac\) so, first find a,b,c then calculate D

OpenStudy (anonymous):

a= x^2 b= 12x C= 12

hartnn (hartnn):

a,b,c are all constants so, comparing ax^2 with x^2 gives a =1 bx with -12 x gives b =-12 got this ??

OpenStudy (anonymous):

144x-4*x^2*12 4*1*12= 48 144-48=96?

hartnn (hartnn):

but yes, \(D = 144-48 = 96\) is correct :)

OpenStudy (anonymous):

How do i put that in a graph tho?

hartnn (hartnn):

*you don't have to graph it* so, if D is positive and is NOT a perfect square, then you will have 2 positive roots. which means it will have 2 points common with x axis

OpenStudy (anonymous):

so its D

hartnn (hartnn):

since your graph opens upwards, and it intersects x axis at 2 points, its vertex will be below x axis

hartnn (hartnn):

not D, no

OpenStudy (anonymous):

Oh ok gotcha so its C then sense its below the axis :)

hartnn (hartnn):

yes, its C :)

OpenStudy (anonymous):

Thanks!! @hartnn